# Empirical validation of the polarization transition in a double-random field model of elections

### v3

@@ Abstract
A central challenge in complex adaptive systems such as societies is the variety of time-dependent interactions, co-evolution, herding, and anticipation of others’ actions [39]. While some aspects are easy to model, calibration and validation remain difficult, requiring great care to ensure testable models. Here, we address opinion formation with multiple interaction types, as in political elections. Voters exchange views within social networks of family, friends, and colleagues, while also following political campaigns—typically only those of their preferred party. We model these two processes through Ising interactions for homophily and a bimodal random field for campaign influence. Together they form a Random Field Ising Model (RFIM), extending spin-spin interactions with site-dependent random external fields.
Despite its simplicity, the RFIM captures rich behavior such as quenched disorder and complex phase diagrams [40, 41, 42]. Variants with bimodal random fields, where the field takes two values [43, 44, 45], exhibit tri-critical points marking the transition from second- to first-order phase changes [46, 47]. In the context of an election campaign, the RFIM represents a bipartisan electorate where each voter holds a binary preference. Voters occupy a social (friendship) network and are randomly assigned one of two field values: the sign encodes campaign affiliation, while the magnitude reflects campaign strength, with spending serving as a proxy. A schematic illustration is shown in End Matter, Fig. 3. RFIM approaches have long been used in sociophysics [48, 49], specifically to illustrate qualitative campaign effects [38], but have never been calibrated to an Ising-type model to with data on campaign spending.
@@
We focus on the effects of increasing campaign intensity. At low spending, opinions are shaped mainly by homophily, but as campaign influence grows, voter preferences are increasingly influenced by campaign messaging. Key questions are: when does campaign alignment outweigh homophilic similarity, how does polarization evolve at this point, and how does this transition affect social tension and election outcomes? A strength of our model is that it can be calibrated and tested on empirical data; we use particularly the US House elections between 1980 and 2020. Calibration allows us to infer the “temperature” and the critical spending threshold above which campaign polarization rises sharply, and to track how many races exceeded this threshold over four decades. To our knowledge, this is the first time where thermodynamic parameters and a critical spending threshold are directly extracted from historical election data and used to predict levels of polarization in society.
@@
To solve the model, we apply two approximations: the configuration model and a mean-field approximation (see Supplemental Material). Denoting the average magnetization by $m=\langle s_{i}\rangle$, we arrive at the mean-field Hamiltonian $H^{MF}(s_{1},\dots,s_{N})=-\sum_{i}(\tilde{J}m+h_{i})s_{i}$, where $\tilde{J}=J\langle k\rangle$. The equilibrium distribution is therefore $p(s|h^{\pm})=\exp\left[-\beta(\tilde{J}m\pm h^{\pm})s\right]/Z^{\pm}$, where $\beta=(kT)^{-1}$ is the inverse temperature (for the rest of the paper, we set $k=1$), and $Z^{\pm}=2\cosh\left[\beta(\tilde{J}m\pm h^{\pm})\right]$ is the partition function. Here, the temperature $T$ represents social volatility—the willingness of individuals to adopt new opinions, even if this increases social stress.
@@
Setting $m=0$ yields
@@
Critical parameters. — We first summarize the known results for the symmetric case $p=\tfrac{1}{2}$ and $h^{+}=h^{-}\equiv h$. We set $\tilde{J}=1$. As shown in [47] and in the Supplemental Material, the model exhibits a continuous crossover for $T>1$. For $T<1$, it undergoes a second-order transition at $h_{c}=T\,\textrm{arctanh}\left(\sqrt{1-T}\right)$. A first-order transition occurs for lower temperatures, with tricritical point $T^{*}=\tfrac{2}{3}$ and $h^{*}=\tfrac{2}{3}\,\textrm{arctanh}\left(\tfrac{1}{\sqrt{3}}\right)\approx 0.439$.
@@
Figure 1: Phase diagrams of the election model. Phase diagram for magnetization, $m$ (a-d), and polarization, $\pi$, (e-h) in the $(h^{+},h^{-})$ plane for temperature $T=1$ (a,c,e,g), and for $T=0.75$ (b,d,f,h), and a prior probability, $p=0.5$ (a,b,e,f) and $p=0.6$ (c,d,g,h). The black dashed line shows $m=0$. The purple point marks the maximal point of the hysteresis, as derived in the main text. For $T=1$, the expected behavior is that the magnetization is directly affected by the relative strength of the two fields, affected by pp. For $T=0.75$, we observe a more interesting behavior of the phase diagram. For the case of low field strength, we observe hysteresis (striped region). In both cases, the campaign polarization, $\pi$, starts to increase rapidly when both field strengths are above a critical value, $h_{c}$ (red).
@@
Figure 2: Estimation of model parameters for US House of Representatives. We compare campaign spending and election results for 6357 of 9135 races between 1980–2020, focusing on close races ($p=0.5\pm 0.05$). (a) For $T\geq 1$, the classification model (see Supplemental Material) predicts the higher-spending candidate wins. (b) For $T<1$, it predicts an incumbency region (yellow) where incumbents win despite lower spending. (c) Optimal parameters $T$ and $h_{c}$ are estimated by maximizing classification accuracy across all 6357 races, yielding $T^{=}0.922$ and $h_{c}=\mathdollar 1.83$M. Cases where incumbents win with lower spending are highlighted, with additional black borders for points in the hysteresis region. The spending diagram is truncated to show the incumbency region (yellow). (d) Accuracy across $T$ is shown, with the maximum marked by a red star. The inset shows a McNemar contingency table comparing the optimal model ($T=T^{*}$) to the null model ($T=1$). The McNemar test gives $p<0.0001$, indicating significantly better performance of the optimal model.
@@
We first consider the symmetric case $p\approx 0.5$. For $p=0.5$, the classification model predicts the same outcome for all $T\geq 1$: the higher-spending candidate wins (Fig. 2(a)). For $T<1$, a hysteresis region emerges (yellow) where incumbency dominates. This region centers near $h^{DEM}=h^{REP}$, with its shape depending on $T$ and $h_{c}$ (Fig. 2(b)). By fitting the empirical boundaries of this region, we infer $T$ and $h_{c}$. Optimal parameters are estimated by maximizing classification accuracy over $T$ and $h_{c}$, yielding $T^{*}=0.922$ and $h^{*}=\mathdollar 1.83M$ (Fig. 2(c)). The figure highlights all closely contested races ($p=0.5\pm 0.05$), marking cases where incumbents win despite lower spending (black-bordered dots).
In the End Matter, we present a similar analysis for races with $p\approx 0.6$, i.e., Republican-leaning races ($p=0.6\pm 0.05$) and $p\approx 0.4$, i.e., Democrat-leaning races. Applied to these two subsets, the parameters shift slightly ($T=0.845$ for Republican-leaning races, $T=0.865$ for Democrat-leaning races, $h^{*}\sim\mathdollar 2M$), but the qualitative behavior — hysteresis and emergence of polarization — remains unchanged.
@@
To further test the robustness of our approach, we present two additinal analyses in the Supplementary Material: in the first analysis, we divide the dataset in four decades and estimate the parameters for each decade separetely. We observe that while the temperature slightly decreases over time, the critical spending slightly increases. Second, we compare our approach to a machine learning approach based Support vector machine classification model. We show that the SVM can estimate the incumbency region (with slightly lower test accuracy), however, it cannot provide additional interpretion of the model, as the presence of polarization region.
Discussion. — We introduced a simple election model combining two mechanisms influencing voter decisions: homophily, i.e., interactions with family, friends, and close contacts, and political campaign influence. Despite its simplicity, the model can be calibrated with real-world data (US House elections) and reproduces rich behavior absent in earlier work. The fact that voters typically follow only one campaign leads to a phase transition: campaign polarization rises sharply once both parties exceed a critical spending threshold. In this regime, most outcomes are 50:50 when $p\approx 0.5$ (swing states), regardless of the detailed spending. For biased states ($p\neq 0.5$), the model allows us to estimate the challenger’s minimum spending above which 50:50 outcomes become possible.
@@
Acknowledgments. — We thank Mirta Galesic and Henrik Olsson for helpful discussions. We acknowledge support from the Austrian Science Fund (FWF) project No. P 34994 (grant-DOI: 10.55776/P34994), the Austrian Research Promotion Agency (FFG) as part of the ESSENCSE project (873927), and the Austrian Federal Ministry for Climate Action, Environment, Mobility, Innovation & Technology (BMK) as part of the CSH Postdoc Programme (GZ 2023-0.841.266).
@@
Detailed derivation of the self-consistency equation. We start with the Hamiltonian
$H(s_{1},\dots,s_{N})=-J\sum_{i<j}A_{ij}s_{i}s_{j}-\sum_{i}h_{i}s_{i}\,$,
In order to decouple the Hamiltonian into the Hamiltonian for a single individual, we use the two approximations. The first one is the configuration model approximation, where we assume that the adjacency matrix of a random network can be approximated $A_{ij}\approx\frac{k_{i}k_{j}}{Nk}$ where $k_{i}$ is the degree (i.e., connectivity) of the node $i$ and $\langle k\rangle$ is the average degree. The second approximation we use is the mean-field approximation. Expressing the spin in terms of its average value as $s_{i}=\langle s_{i}\rangle+\delta s_{i}$ enables us to omit the term quadratic in fluctuations. By denoting the average magnetization as $m=\langle s_{i}\rangle$, we can rewrite the Hamiltonian as
$H(s_{1},\dots,s_{N})$ $\approx$ $\frac{Jm^{2}N\langle k\rangle}{2}-J\sum_{i}\langle k\rangle ms_{i}-\sum_{i}h_{i}s_{i}$
The first term can be omitted from the Hamiltonian since it is just an additive constant to the energy, which will be canceled out once the equilibrium distribution is calculated. Thus, the mean-field Hamiltonian can be expressed as
@@
where $\beta=\frac{1}{kT}$ is the inverse temperature, and
@@
$m^{\pm}=\langle s\rangle^{\pm}=\sum_{s=\pm 1}sp(s|h^{\pm})=\tanh\left(\beta(\tilde{J}m\pm h^{\pm})\right)$
. The total magnetization can be expressed as
@@
Derivation of the critical curve and tricritical point for symmetric case. — Let us now focus on the symmetric case, i.e., when $p=\frac{1}{2}$ and $h^{+}=h^{-}\equiv h$. We investigate how the phase diagram depends on the strength of the field and the temperature. Without loss of generality, we consider that $\tilde{J}=1$. To this end, we expand the right-hand side of the self-consistency equation around $m=0$ and get
@@
$m^{\pm}=\pm\frac{\sqrt{3(1-T-\xi^{2}})}{\sqrt{\frac{T^{2}+3\xi^{4}-4\xi^{2}}{T^{2}}}}$
@@
$h_{c}=T\textrm{arctanh}(\sqrt{1-T})$
@@
$T^{*}=\frac{2}{3}\quad h^{*}=\frac{2}{3}\textrm{arctanh}\left(\frac{1}{\sqrt{3}}\right)\approx 0.439$
Derivation of the critical curve and tricritical point for the asymmetric case. Let us now focus on the general case when $p\neq\frac{1}{2}$. We take the condition for $m=0$, which is
@@
around $m=0$ while keeping the dependence between $h^{+}$ and $h^{-}$ determined from the condition on $m=0$ above.
Let us use the Taylor expansion of $\tanh(\beta(m+h))$ which is
$\tanh(\beta(m\pm h^{\pm}))=\pm\tanh(\beta h^{\pm})$
$+\beta m\left(1-\tanh^{2}(\beta h^{\pm})\right)$
$\mp\beta^{2}m^{2}\tanh(\beta h^{\pm})(1-\tanh^{2}(\beta h^{\pm}))+\mathcal{O}(m^{3})$
@@
$m=p\left[\xi^{+}+m\frac{1-(\xi^{+})^{2}}{T}-m^{2}\frac{\xi^{+}(1-(\xi^{+})^{2})}{T^{2}}\right]$
$+(1-p)\left[-\xi^{-}+m\frac{1-(\xi^{-})^{2}}{T}+m^{2}\frac{\xi^{-}(1-(\xi^{-})^{2})}{T^{2}}\right]+\mathcal{O}(m^{3})$
Since $p\xi^{+}-(1-p)\xi^{-}=0$, the constant term is zero; therefore, we can rearrange the terms as
$\left[\frac{T-p(1-(\xi^{+})^{2})-(1-p)(1-(\xi^{-})^{2})}{T}\right]m$
$+\left[\frac{p\xi^{+}(1-(\xi^{+})^{2})-(1-p)\xi^{-}(1-(\xi^{-})^{2})}{T^{2}}\right]m^{2}=0$
@@
$T=p(1-(\xi^{+}_{c})^{2})+(1-p)(1-\left(\frac{p}{1-p}\right)^{2}(\xi^{+}_{c})^{2})$
$=1-\frac{p}{1-p}(\xi^{+}_{c})^{2}\,.$
@@
Classification model. Here, we describe the classification models used in the main text. The classification model is directly based on the results of the Random Field Ising model. Without loss of generality, we assign $h^{+}=h^{REP}$ as the spending of the Republican party candidate, $h^{-}=h^{DEM}$ as the spending of the Democratic party candidate. The prediction of the classification model, based on the magnetization $m$ goes as follows:
@@
  • If $(h^{DEM},h^{REP})$ lie outside of the hysteresis region, or if there is an open seat (i.e., the incumbent does not candidate), a Republican wins if $m>0$, and a Democrat wins if $m<0$.
@@
Model accuracy. In order to measure the performance of the classification model, we use the model accuracy. The confusion table between the predicted classification and the actual classification (here, the prediction is that a Republican candidate wins an election) is then defined as
@@
positive True positive (TP) False negative (FN)
negative False positive (FP) True negative (TN)
The accuracy is defined as
$ACC=\frac{n_{TP}+n_{TN}}{n_{TP}+n_{TN}+n_{FP}+n_{FN}}$
McNemar test. The McNemar test is used to demonstrate whether one of the two classification models used on a given data set is better than the other. For each observation, a classification model gives a predicted classification, which is compared with the actual classification. For example, in the election races, the classification model predicts the winner of the election based on the campaign spending and incumbency (see the second above), which is then compared with the actual election result.
For two classification models $M_{1}$$M_{2}$, the contingency table between correctly and incorrectly classified observations can be written as follows:

### v4

@@ Abstract
A central challenge in complex adaptive systems such as societies is the variety of time-dependent interactions, co-evolution, herding, and anticipation of others’ actions [39]. While some aspects are easy to model, calibration and validation remain difficult, requiring great care to ensure testable models. Here, we address opinion formation with multiple interaction types, as in political elections. Voters exchange views within social networks of family, friends, and colleagues, while also following political campaigns—typically only those of their preferred party. We model these two processes through Ising interactions for homophily and a bimodal random field for campaign influence. Together, they form a Random Field Ising Model (RFIM), extending spin-spin interactions with site-dependent random external fields. An equivalent mean-field description can be derived from a master-equation of stochastic opinion switching with preference and adaptation terms, as originally introduced in the context of sociodynamics [40, 41].
Despite its simplicity, the RFIM captures rich behavior such as quenched disorder and complex phase diagrams [42, 43, 44]. Variants with bimodal random fields, where the field takes two values [45, 46, 47], exhibit tri-critical points marking the transition from second- to first-order phase changes [48, 49]. In the context of an election campaign, the RFIM represents a bipartisan electorate where each voter holds a binary preference. Voters occupy a social (friendship) network and are randomly assigned one of two field values: the sign encodes campaign affiliation, while the magnitude reflects campaign strength, with spending serving as a proxy. A schematic illustration is shown in End Matter, Fig. 3. RFIM approaches have long been used in sociophysics [50, 51], specifically to illustrate qualitative campaign effects [38], but have never been calibrated to an Ising-type model with data on campaign spending. The model employs concepts such as temperature and external fields, which should be understood as effective parameters that summarize complex social, cognitive, and economic processes and not as literally physical quantities.
@@
We focus on the effects of increasing campaign intensity. At low spending, opinions are shaped mainly by homophily, but as campaign influence grows, voter preferences are increasingly influenced by campaign messaging. Key questions are: when does campaign alignment outweigh homophilic similarity, how does polarization evolve at this point, and how does this transition affect social tension and election outcomes? A strength of our model is that it can be calibrated and tested on empirical data; we use the US House elections between 1980 and 2020, particularly. Calibration allows us to infer the “temperature” and the critical spending threshold above which campaign polarization rises sharply, and to track how many races exceeded this threshold over four decades. To our knowledge, this is the first time that thermodynamic parameters and a critical spending threshold are directly extracted from historical election data and used to predict levels of polarization in society.
@@
To solve the model, we apply two approximations: the configuration model and a mean-field approximation (see Supplemental Material). Denoting the average magnetization by $m=\langle s_{i}\rangle$, we arrive at the mean-field Hamiltonian $H^{MF}(s_{1},\dots,s_{N})=-\sum_{i}(\tilde{J}m+h_{i})s_{i}$, where $\tilde{J}=J\langle k\rangle$ and $\langle k\rangle$ is the average node degree. The equilibrium distribution is therefore $p(s|h^{\pm})=\exp\left[-\beta(\tilde{J}m\pm h^{\pm})s\right]/Z^{\pm}$, where $\beta=(kT)^{-1}$ is the inverse temperature (for the rest of the paper, we set $k=1$), and $Z^{\pm}=2\cosh\left[\beta(\tilde{J}m\pm h^{\pm})\right]$ is the partition function. Here, the temperature $T$ represents social volatility—the willingness of individuals to adopt new opinions, even if this increases social stress.
@@
Equation (3) can be alternatively derived from a master-equation of stochastic opinion switching with preference and adaptation terms; the explicit derivation and parameter mapping are given in the Supplemental Material. Setting $m=0$ yields
@@
Critical parameters. — We first summarize the known results for the symmetric case $p=\tfrac{1}{2}$ and $h^{+}=h^{-}\equiv h$. We set $\tilde{J}=1$. As shown in [49] and in the Supplemental Material, the model exhibits a continuous crossover for $T>1$. For $T<1$, it undergoes a second-order transition at $h_{c}=T\,\textrm{arctanh}\left(\sqrt{1-T}\right)$. A first-order transition occurs for lower temperatures, with tricritical point $T_{t}=\tfrac{2}{3}$ and $h_{c}=\tfrac{2}{3}\,\textrm{arctanh}\left(\tfrac{1}{\sqrt{3}}\right)\approx 0.439$.
@@
Figure 1: Phase diagrams of the election model. Phase diagram for magnetization, $m$ (a-d), and polarization, $\pi$, (e-h) in the $(h^{+},h^{-})$ plane for temperature $T=1$ (a,c,e,g), and for $T=0.75$ (b,d,f,h), and a prior probability, $p=0.5$ (a,b,e,f) and $p=0.6$ (c,d,g,h). The black dashed line shows $m=0$. The purple point marks the maximal point of the hysteresis, as derived in the main text. For $T=1$, the expected behavior is that the magnetization is directly affected by the relative strength of the two fields, affected by pp. For $T=0.75$, we observe a more interesting behavior of the phase diagram. For the case of low field strength, we observe hysteresis (striped region). In both cases, the campaign polarization, $\pi$, increases rapidly when both field strengths exceed a critical value, $h_{c}$ (red).
@@
Figure 2: Estimation of model parameters for US House of Representatives. We compare campaign spending and election results for 6357 of 9135 races between 1980–2020, focusing on close races ($p=0.5\pm 0.05$). (a) For $T\geq 1$, the classification model (see Supplemental Material) predicts the higher-spending candidate wins. (b) For $T<1$, it predicts an incumbency region (yellow) where incumbents win despite lower spending. (c) Optimal parameters $T$ and $h_{c}$ are estimated by maximizing classification accuracy across all 6357 races, yielding $T^{*}=0.922$ and $h_{c}=\mathdollar 1.83$M. Cases where incumbents win with lower spending are highlighted, with additional black borders for points in the hysteresis region. The spending diagram is truncated to show the incumbency region (yellow). (d) Accuracy across $T$ is shown, with the maximum marked by a red star. The inset shows a McNemar contingency table comparing the optimal model ($T=T^{*}$) to the null model ($T=1$). The McNemar test gives $p<0.0001$, indicating significantly better performance of the optimal model.
@@
We first consider the symmetric case $p\approx 0.5$. For $p=0.5$, the classification model predicts the same outcome for all $T\geq 1$: the higher-spending candidate wins (Fig. 2(a)). For $T<1$, a hysteresis region emerges (yellow) where incumbency dominates. This region centers near $h^{DEM}=h^{REP}$, with its shape depending on $T$ and $h_{c}$ (Fig. 2(b)). By fitting the empirical boundaries of this region, we infer $T$ and $h_{c}$. Optimal parameters are estimated by maximizing classification accuracy over $T$ and $h_{c}$, yielding $T^{*}=0.922$ and $h_{c}=\mathdollar 1.83M$ (Fig. 2(c)). The figure highlights all closely contested races ($p=0.5\pm 0.05$), marking cases where incumbents win despite lower spending (black-bordered dots).
In the End Matter, we present a similar analysis for races with $p\approx 0.6$, i.e., Republican-leaning races ($p=0.6\pm 0.05$, Fig. 5) and $p\approx 0.4$, i.e., Democrat-leaning races ($p=0.4\pm 0.05$, Fig. 6). Applied to these two subsets, the parameters shift slightly ($T^{*}=0.845$ for Republican-leaning races, $T^{*}=0.865$ for Democrat-leaning races, $h_{c}\sim\mathdollar 2M$), but the qualitative behavior — hysteresis and emergence of polarization — remains unchanged.
@@
To further test the robustness of our approach, we present two additional analyses in the Supplemental Material: in the first analysis, we divide the dataset into four decades and estimate the parameters for each decade separately. We observe that while the temperature slightly decreases over time, the critical spending slightly increases. Second, we compare our approach to a machine learning approach based on a support vector machine classification model. We show that the SVM can estimate the incumbency region (with slightly lower test accuracy); however, it cannot provide additional interpretation of the model, such as the presence of a polarization region.
Discussion. — We introduced a simple election model combining two mechanisms influencing voter decisions: homophily, i.e., interactions with family, friends, and close contacts, and political campaign influence. Despite its simplicity, the model can be calibrated with real-world data (US House elections) and reproduces rich behavior absent in earlier work. The fact that voters typically follow only one campaign leads to a phase transition: campaign polarization rises sharply once both parties exceed a critical spending threshold. In this regime, most outcomes are 50:50 when $p\approx 0.5$ (swing states), regardless of the detailed spending. For biased states ($p\not\approx 0.5$), the model allows us to estimate the challenger’s minimum spending above which 50:50 outcomes become possible.
@@
Acknowledgments. — We thank Mirta Galesic and Henrik Olsson for helpful discussions. We also thank three anonymous referees for their constructive comments. We acknowledge support from the Austrian Science Fund (FWF) under Grants No. 10.55776/P34994 and EFP5 ReMASS, funding from the Austrian Federal Ministry for Climate Action, Environment, Energy, Mobility, Innovation, and Technology under GZ 2023-0.841.266, through the Postdoc Program for Complexity Science and Data Competence.
Data availability statement. — he data that support the findings of this article are openly available [59, 60].
@@
Detailed derivation of the self-consistency equation
We start with the Hamiltonian
$H(s_{1},\dots,s_{N})=-J\sum_{i<j}A_{ij}s_{i}s_{j}-\sum_{i}h_{i}s_{i}\,.$
In order to decouple the Hamiltonian into the Hamiltonian for a single individual, we use the two approximations. The first one is the configuration model approximation, where we assume that the adjacency matrix of a random network can be approximated as $A_{ij}\approx\frac{k_{i}k_{j}}{N\langle k\rangle}$, where $k_{i}$ is the degree (i.e., connectivity) of the node $i$ and $\langle k\rangle$ is the average degree. The second approximation we use is the mean-field approximation. Expressing the spin in terms of its average value as $s_{i}=\langle s_{i}\rangle+\delta s_{i}$ enables us to omit the term quadratic in fluctuations. By denoting the average magnetization as $m=\langle s_{i}\rangle$, we can rewrite the Hamiltonian as
$H(s_{1},\dots,s_{N})$ $\approx$ $\frac{Jm^{2}N\langle k\rangle}{2}-J\sum_{i}\langle k\rangle\,ms_{i}-\sum_{i}h_{i}s_{i}\,.$
The first term can be omitted from the Hamiltonian, as it is an additive constant of the energy and will be canceled when the equilibrium distribution is calculated. Thus, the mean-field Hamiltonian can be expressed as
@@
where $\beta=\frac{1}{kT}$ is the inverse temperature (we set $k=1$), and
@@
$m^{\pm}=\langle s\rangle^{\pm}=\sum_{s=\pm 1}sp(s|h^{\pm})=\tanh\left(\beta(\tilde{J}m\pm h^{\pm})\right)\,.$
The total magnetization can be expressed as
@@
Derivation of the critical curve and tricritical point for symmetric case
Let us now focus on the symmetric case, i.e., when $p=\frac{1}{2}$ and $h^{+}=h^{-}\equiv h$. We investigate how the phase diagram depends on the strength of the field and the temperature. Without loss of generality, we consider that $\tilde{J}=1$. To this end, we expand the right-hand side of the self-consistency equation around $m=0$ and get
@@
$m^{\pm}=\pm\frac{\sqrt{3(1-T-\xi^{2}})}{\sqrt{\frac{T^{2}+3\xi^{4}-4\xi^{2}}{T^{2}}}}\,.$
@@
$h_{c}=T\textrm{arctanh}(\sqrt{1-T})\,.$
@@
$T^{*}=\frac{2}{3}\,,\qquad h^{*}=\frac{2}{3}\,\textrm{arctanh}\left(\frac{1}{\sqrt{3}}\right)\approx 0.439\,.$
Derivation of the critical curve and tricritical point for the asymmetric case
Let us now focus on the general case when $p\neq\frac{1}{2}$. We take the condition for $m=0$, which is
@@
around $m=0$ while keeping the dependence between $h^{+}$ and $h^{-}$ determined from the condition on $m=0$ above. We use the Taylor expansion of $\tanh(\beta(m+h))$ which is
$\tanh(\beta(m\pm h^{\pm}))=\pm\tanh(\beta h^{\pm})+\beta m\left(1-\tanh^{2}(\beta h^{\pm})\right)\mp\beta^{2}m^{2}\tanh(\beta h^{\pm})(1-\tanh^{2}(\beta h^{\pm}))+\mathcal{O}(m^{3})\,.$
@@
$m=p\left[\xi^{+}+m\frac{1-(\xi^{+})^{2}}{T}-m^{2}\frac{\xi^{+}(1-(\xi^{+})^{2})}{T^{2}}\right]+(1-p)\left[-\xi^{-}+m\frac{1-(\xi^{-})^{2}}{T}+m^{2}\frac{\xi^{-}(1-(\xi^{-})^{2})}{T^{2}}\right]+\mathcal{O}(m^{3})\,.$
Since $p\,\xi^{+}-(1-p)\,\xi^{-}=0$, the constant term is zero. Therefore, we can rearrange the terms as
$\left[\frac{T-p(1-(\xi^{+})^{2})-(1-p)(1-(\xi^{-})^{2})}{T}\right]m+\left[\frac{p\xi^{+}(1-(\xi^{+})^{2})-(1-p)\xi^{-}(1-(\xi^{-})^{2})}{T^{2}}\right]m^{2}=0\,.$
@@
$T=p(1-(\xi^{+}_{c})^{2})+(1-p)\left[1-\left(\frac{p}{1-p}\right)^{2}(\xi^{+}_{c})^{2}\right]=1-\frac{p}{1-p}(\xi^{+}_{c})^{2}\,.$
@@
Alternative derivation of the self-consistency equation from the Weidlich master-equation model
In this section, we show that the mean-field description used in the main text can alternatively be obtained from a master-equation approach to opinion dynamics, following the sociodynamics framework introduced by Weidlich and Haag. This demonstrates that the polarization transition does not rely on the Hamiltonian formulation with spatial interactions, but emerges generically from stochastic opinion switching driven by individual preferences and social adaptation.
Let us consider a population of $N$ voters with binary opinions $s_{i}\in\pm 1$. Let us define the number of voters with a positive opinion as $n$, and the fraction of voters with opinion $s=+1$ as $x=n/N\in[0,1]$. The natural connection to the magnetization defined in the main paper is m=2x1m=2x-1. The system undergoes a Markovian evolution described by a master equation, fully characterized by the transition rates $W_{+1\,\rightarrow\,-1}(n)\equiv W_{+-}(n)$ and $W_{-1\,\rightarrow\,+1}(n)\equiv W_{-+}(n)$. The master equation can therefore be expressed as
$\frac{\mathrm{d}P(n)}{\mathrm{d}t}$ $=$ $(N-n+1)W_{-+}(n-1)P(n-1,t)$
$+$ $(n+1)W_{+-}(n+1)P(n+1,t)$
- $[(N-n)W_{-+}(n)+nW_{+-}(n)]P(n,t)\,.$
In the classic formulation of sociodynamics, the individual opinion changes are governed by two conceptually distinct mechanisms. First, individuals possess intrinsic preferences that are independent of the current social configuration and may reflect long-standing inclinations, prior beliefs, or external information. In the election setting, this corresponds to the election campaign they follow
Second, individuals exhibit an adaptive response to the prevailing opinion, whereby the propensity to adopt a given opinion increases with its share of the population, reflecting conformity or social pressure. This mechanism is somewhat similar to the homophily, although it does not necessarily require the assumption on the spatial distribution of interactions. These preference and adaptation tendencies act simultaneously at the individual level and may have independent strengths. The particular choice that is widely used in sociodynamics literature is
$W_{-+}(n)$ $=$ $\nu\exp(\theta+Kx)\,$,
$W_{+-}(n)$ $=$ $\nu\exp(-(\theta+Kx))\,$,
where $\theta$ quantifies the former mechanism, while $Kx$ the latter one. In our setting, $\theta$ represents the strength of the election campaign. Thus, in our scenario, we divide the population into two subpopulations, one following the campaign of $s=+1$ of size $N^{+}$ and the other following the campaign of s=1s=-1 of size $N^{-}$. We denote the number of individuals following the first campaign with a positive opinion as $n^{+}$ and analogously $n^{-}$. The campaign intensities are $\theta^{+}$ and $\theta^{-}$, and the transition rates for gives subpopulations are therefore
$W_{-+}^{+}(n^{+})$ $=$ $\nu\exp(\theta^{+}+Kx)\,$,
$W_{+-}^{+}(n^{+})$ $=$ $\nu\exp(-(\theta^{+}+Kx))\,$,
$W_{-+}^{-}(n^{-})$ $=$ $\nu\exp(-\theta^{-}+Kx)\,$,
$W_{+-}^{-}(n^{-})$ $=$ $\nu\exp(-(-\theta^{-}+Kx))\,.$
Here, individuals pursue their own campaigns but still seek to align with the majority of the population, regardless of the campaign. This is reflected by the term $Kx$. By calculating the first moment of the probability distribution from the master equation for each population in the case of $N\gg 1$, we obtain
$\dot{x}^{+}$ $=$ $(1-x^{+})\nu e^{\theta^{+}+Kx}-x^{+}\nu e^{-(\theta^{+}+Kx)}\,$,
$\dot{x}^{-}$ $=$ $(1-x^{-})\nu e^{-\theta^{-}+Kx}-x^{-}\nu e^{-(-\theta^{-}+Kx)}\,.$
At the stationary point, we obtain, after a straightforward calculation
$x^{\pm}=\frac{1}{2}\left[1+\tanh\!\big(Kx\pm\theta^{\pm}\big)\right]\,.$
By defining $p=N^{+}/N$, we can rewrite normalization as $x=px^{+}+(1-p)x^{-}$, and therefore we obtain
$x=p\,\frac{1}{2}\left[1+\tanh(Kx+\theta^{+})\right]+(1-p)\,\frac{1}{2}\left[1+\tanh(Kx-\theta^{-})\right]\,.$
Using the relation m=2x1m=2x-1, the term $Kx$ can be rewritten as $\frac{Km}{2}+\frac{K}{2}$ where the constant contribution is absorbed into a redefinition of the effective fields. Thus, by choosing the following transformation
$m$ $=$ $2x-1\,$,
$2\beta\tilde{J}$ $=$ $K\,$,
$\beta h^{+}$ $=$ $\theta^{+}+\frac{K}{2}\,$,
$\beta h^{-}$ $=$ $\theta^{-}-\frac{K}{2}\,$,
we obtain the equation for $m$
$m=p\tanh\left(\beta(\tilde{J}m+h^{+})\right)+(1-p)\tanh\left(\beta(\tilde{J}m-h^{-})\right)$
which is exactly the self-consistency equation from the main text.
Classification model
Here, we describe the classification models used in the main text. The classification model is directly based on the results of the Random Field Ising model. Without loss of generality, we assign $h^{+}=h^{REP}$ as the spending of the Republican party candidate, $h^{-}=h^{DEM}$ as the spending of the Democratic party candidate. The prediction of the classification model, based on the magnetization $m$ goes as follows:
@@
  • If $(h^{DEM},h^{REP})$ lie outside of the hysteresis region, or if there is an open seat (i.e., the incumbent does not run as a candidate), a Republican wins if $m>0$, and a Democrat wins if $m<0$.
@@
Model accuracy
In order to measure the performance of the classification model, we use the model accuracy. The confusion table between the predicted classification and the actual classification (here, the prediction is that a Republican candidate wins an election) is then defined in Tab. 1. The accuracy is defined as
$ACC=\frac{n_{TP}+n_{TN}}{n_{TP}+n_{TN}+n_{FP}+n_{FN}}\,.$
@@
positive True positive ($n_{TP}$) False negative ($n_{FN}$)
negative False positive ($n_{FP}$) True negative ($n_{TN}$)
Table 1: Confusion table of the classification model.
McNemar test
The McNemar test is used to demonstrate whether one of the two classification models used on a given data set is better than the other. For each observation, a classification model gives a predicted classification, which is compared with the actual classification. For example, in the election races, the classification model predicts the winner of the election based on the campaign spending and incumbency (see the second above), which is then compared with the actual election result. For two classification models $M_{1}$$M_{2}$, the contingency table between correctly and incorrectly classified observations can be written as shown in Tab. 2.